# The Impact of Surface Co-location and Eye-tracking on Mixed Reality Typing

### v2

@@ 5. Conclusion
Experiment 1 showed that a midair keyboard had significantly higher entry rates and was preferred over keyboards co-located on a surface. Entry rates were 12.2, 9.1, and 7.8 WPM in the Midair, Table, and Wall conditions respectively. Questionnaire results also showed that our midair keyboard, which was arranged such that participants could rest their arms on a table, was preferred and not particularly tiring compared to the wall and table co-located keyboards. Participant feedback was generally positive about the tactile feedback and ergonomics of typing on a virtual keyboard located on a table. However, at present, the hand-tracking of a current commodity headset (HoloLens 2) probably need improvement to better support accurate deterministic typing near surfaces.
In Experiment 2, we further explored midair typing by implementing typing with ten fingers and using eye-tracking to help eliminate unintentional key presses and comparing it to index-finger typing. While we found eye-tracking reduced the need to correct errors, ten-finger typing in midair was still slower than typing with only the index fingers. Entry rates were 10.9, 9.4, 7.1, and 7.4 WPM in the Index, IndexEye, Ten, and TenEye conditions respectively. In situations where midair ten-finger typing is desired, our results suggest that the addition of eye-tracking can significantly reduce backspace use without negatively impacting either entry or final error rates. Questionnaire results also showed no significant negative impact of the addition of eye-tracking.
@@ Acknowledgements.
- Adhikary and Vertanen (2021) Jiban Adhikary and Keith Vertanen. 2021. Text Entry in Virtual Environments using Speech and a Midair Keyboard. IEEE Transactions on Visualization and Computer Graphics 27, 5 (2021), 2648–2658. https://doi.org/10.1109/TVCG.2021.3067776

### v3

@@ 6. Conclusion
Experiment 1 showed that a midair keyboard had significantly higher entry rates and was preferred over keyboards positioned on a surface. Entry rates were 12.2, 9.1, and 7.8 WPM in the Midair, Table, and Wall conditions respectively. Questionnaire results also showed that our midair keyboard, which was arranged such that participants could rest their arms on a table, was preferred and not particularly tiring compared to the keyboards positioned on a wall or table. Participant feedback was generally positive about the tactile feedback and ergonomics of typing on a virtual keyboard located on a table. However, at present, the hand-tracking of a current commodity headset (HoloLens 2) needs improvement to better support accurate deterministic interactions near surfaces.
In Experiment 2, we further explored midair typing comparing typing with ten fingers versus typing with index fingers. We tested using eye-tracking to help filter unintentional key presses. While we found eye-tracking reduced error corrections in ten-finger typing, it was still slower than index-finger typing. Entry rates were 10.9, 9.4, 7.1, and 7.4 WPM in the Index, IndexEye, Ten, and TenEye conditions respectively. In situations where midair ten-finger typing is desired, our results suggest eye-tracking filtering can reduce backspace use without negatively impacting users’ entry or error rates. Subjective feedback also showed participants felt eye-tracking helped them type more accurately.
@@ Acknowledgements.
- Adhikary and Vertanen (2021a) Jiban Adhikary and Keith Vertanen. 2021a. Text Entry in Virtual Environments using Speech and a Midair Keyboard. IEEE Transactions on Visualization and Computer Graphics 27, 5 (2021), 2648–2658. https://doi.org/10.1109/TVCG.2021.3067776
- Adhikary and Vertanen (2021b) Jiban Adhikary and Keith Vertanen. 2021b. Typing on Midair Virtual Keyboards: Exploring Visual Designs and Interaction Styles. In Human-Computer Interaction – INTERACT 2021: 18th IFIP TC 13 International Conference, Bari, Italy, August 30 – September 3, 2021, Proceedings, Part IV (Bari, Italy). Springer-Verlag, Berlin, Heidelberg, 132–151. https://doi.org/10.1007/978-3-030-85610-6_9
@@
- Chiossi et al. (2024) Francesco Chiossi, Yassmine El Khaoudi, Changkun Ou, Ludwig Sidenmark, Abdelrahman Zaky, Tiare Feuchtner, and Sven Mayer. 2024. Evaluating Typing Performance in Different Mixed Reality Manifestations using Physiological Features. Proceedings of the ACM on Human-Computer Interaction 8, ISS, Article 542 (Oct. 2024), 30 pages. https://doi.org/10.1145/3698142
@@
- Dudley (2025) John J. Dudley. 2025. Personal communication. Email correspondence.
@@
- Wobbrock et al. (2011) Jacob O Wobbrock, Leah Findlater, Darren Gergle, and James J Higgins. 2011. The aligned rank transform for nonparametric factorial analyses using only anova procedures. In Proceedings of the SIGCHI conference on human factors in computing systems. 143–146.